OpenStudy (jennychan12):

Find d^2y/dx^2 (second derivative) of x(siny) = 3

OpenStudy (jennychan12):

i found that dy/dx = -tany/x

OpenStudy (jennychan12):

so i got d^2y/dx^2 = -tany/x^2

OpenStudy (jennychan12):

is that right?

OpenStudy (anonymous):

not sure with the second derivative

OpenStudy (anonymous):

\[\frac{d^2y}{dx^2} = -\frac{ x\frac{dy}{dx}\sec^2 y - \tan y }{x^2}\]\[\frac{d^2y}{dx^2} = \frac{\tan y \sec^2 y + \tan y}{x^2}\]

OpenStudy (jennychan12):

ok thanks. lemme see what i did wrong.

OpenStudy (jennychan12):

wait. i think d^2y/dx^2 is wrong. shouldn't it be  ? i can do the algebra from here...

OpenStudy (anonymous):

you forget to multiply by -1, and dy/dx = -(tan y)/x, not -(tan y)/x^2

OpenStudy (jennychan12):

ohh i see. sorry i looked at the wrong dy/dx

OpenStudy (anonymous):

Here's sort of what I'm getting \[ \begin{array}{rcl} x(\sin y) &=& 3 \\ \frac{d}{dx}x(\sin y) &=& \frac{d}{dx}3 \\ \sin y + x\frac{dy}{dx}\cos y &=& 0 \\ x\frac{dy}{dx}\cos y &=& -\sin y \\ x\frac{dy}{dx} &=& -\tan y \\ \frac{dy}{dx} &=& \frac{ -\tan y}{x} \\ \frac{d}{dx} \left[ \frac{dy}{dx} \right] &=& \frac{d}{dx} \left[ \frac{- \tan y}{x} \right] \\ \frac{d^2y}{dx^2} &=& \frac{-\sec^2(y)\frac{dy}{dx}x-(-\tan y)}{x^2} \\ \end{array} \]

OpenStudy (anonymous):

Would have to sub in the \(\frac{-\tan y}{x}\) for \(\frac{dy}{dx}\) at the end.

OpenStudy (jennychan12):

yeah i know. thanks. i think i originally took the wrong derivative